OpenStudy (cderaalien):

In this equation: x + 4 = -2x - 2 Is the 'x' variable the same thing as '1x'? and to solve this equation, do I start by subtracting the 'x' from both sides, or -2x from both sides?

OpenStudy (irishboy123):

yes you can sub x or add 2x.

OpenStudy (irishboy123):

hope that wasn't too terse btw!!!! ask if more questions arise..... :p

OpenStudy (cderaalien):

@IrishBoy123 okay lol soo...I did the problem and I got the solution (-6,10) but this solution only works for one equation? What did I do wrong?

OpenStudy (irishboy123):

\(x + 4 = -2x - 2\) \(x \color{red}{+2x}+ 4 = -2x \color{red}{+2x}- 2\) \(3x + 4 = -2\) guess, from your answer, you didn't get that far...??

OpenStudy (cderaalien):

I got far, I just subtracted x from both sides. The two x's on the left side of the equation canceled out, and on the right side I got the answer of -1x (because -2x - x). So, now the equation looked like this: 4 = -1x - 2 So, I added 2 to each side of the equation, the 2's on the right side cancel each other out and then on the left side 2 + 4 = 6. So, now the problem looked like this: 6 = -1x. I, then, divided -1x/-1 & 6/-1. It concluded to, -6 = x.

OpenStudy (cderaalien):

@IrishBoy123, where did I mess up?

OpenStudy (irishboy123):

ok, let go that way instead.... \[x + 4 = -2x - 2\] \[x - \color{red}{x}+ 4 = -2x - \color{red}{x}- 2\] \[4 = - \color{red}{3}x -2\]